OpenStudy (anonymous):

Trig. Identities: Solve from one side only please to prove it cos(3x) = 4 cos^3 x - 3 cos x

OpenStudy (dumbcow):

ok cos(3x) = cos(2x+x) using angle addition identity cos(2x+x) = cos(2x)cos(x) - sin(2x)sin(x) ... using double angle identities cos(2x) = 2cos(x)^2 - 1 sin(2x) = 2sin(x)cos(x) ... = (2cos(x)^2 - 1)*cos(x) - (2sin(x)cos(x))*sin(x) =2cos(x)^3 - cos(x) - 2cos(x)sin(x)^2 ... using trig identity sin(x)^2 = 1-cos(x)^2 ... =2cos(x)^3-cos(x)-2cos(x)(1-cos(x)^2) =4cos(x)^3 - 3cos(x)

OpenStudy (dumbcow):

hope you can follow that

OpenStudy (anonymous):

Yes, that is sooooo helpful! Thanks a bunch, I was really stuck!

OpenStudy (dumbcow):

your welcome, its really just knowing which identies to use you to play around with it a little bit